OpenStudy (anonymous):

x+1/x=3,then how much is

OpenStudy (anonymous):

\[x^7+1/x^7\]

OpenStudy (anonymous):

\[x^2+2+1/x^2=3^2\] \[x^2+1/x^2=7\]

OpenStudy (anonymous):

\[x^3+y^3=(x+y)(x^2-xy+y^2)\] \[x^3+1/x^3=(3)(7)=21\]

OpenStudy (anonymous):

can i make this a sequence \[3,7,21,...\]

OpenStudy (anonymous):

sorry \[(3)(7-1)=18\] \[3,7,18..\]

OpenStudy (anonymous):

\[x^4+1/x^4=49-2=47\]